OpenStudy (vera_ewing):

Given that f(x) = 3x + 1 and g(x) = (4x+2)/3 , solve for g(f(0)).

OpenStudy (gorv):

first find f(0)

OpenStudy (anonymous):

First you find g(x) when x = 0 so that would be: (0+2)/3 = 2/3 Then you evaluate your f(x) with 2/3 as x, I believe.

OpenStudy (gorv):

for finding f(0) plug in x=0 in f(x)

OpenStudy (vera_ewing):

f(x) = 3(0) + 1 f(x)=0+1 f(x)=1

OpenStudy (vera_ewing):

now what?

OpenStudy (anonymous):

Oh god , sorry did that wrong: it's when f(x) has x = 0 so f(x) = 1 and then you plug in 1 for x in g(x)

OpenStudy (vera_ewing):

I got 8 as my answer is that right?

OpenStudy (anonymous):

(4x + 2)/3 (4 +2)/3 6/3 2

OpenStudy (vera_ewing):

Thanks @isagnik ! :)